A particle of mass m is acted upon by a constant power P.The distance travelled by the particle when its velocity increases from v1 to v2 is ?

Grade:12

4 Answers

sushant singh
13 years ago

prankur ,

 what is the question dear...????

Prankur Kataria
13 years ago

We are given a particle of mass m which is acted upon by a constant power P.......What is the distance travelled by the particle when its velocity increases from V1 to V2 ????

Umesh Umesh
6 years ago
Don`t waste your timely reply to this is a good to go to the best of luck for my updated CV and cover letter to this is a good
Rizul
5 years ago
P=FV
P=mav
P=mv . dv/ds . v
P/m.ds =v^2.dv
Integrate both sides under limits
V from v1 to v2
ds from 0 to S
We get
(v^3/3)v1tov2=P/m.s
1/3(v2^3-v1^3)=P/m.S
M/3P(v2^3-v1^3)=S